# DatAasee: A Metadata-Lake as Metadata Catalog for a Virtual Data-Lake

### v1

@@ Abstract
Metadata management for distributed data sources is a long-standing but ever-growing problem. To counter this challenge in a research-data and library-oriented setting, this work constructs a data architecture, derived from the data-lake: the metadata-lake. A proof-of-concept implementation of this proposed metadata system is presented and evaluated as well.
@@ 1 Introduction
In the shadows of data evolving to big data, metadata became big metadata. Because of metadata’s idiosyncrasy of describing other data, metadata management for big data warrants its own discussion. Hence, this inert dependency of metadata on its data needs to be accounted for, and leads to the driving conjecture behind this work and its object of study: existence, detail and accuracy of metadata determines discoverability, use and value of its associated data. To this end, a system for metadata management, which centralizes metadata from distributed data sources, is defined and its implementation described in the following.
A starting point is the abstract data architecture behind such a system: A data-lake is a central repository for the storage of (raw) data, typically together with at least some facility to ingest data (Mathis 2017; Serra 2024). The rationale for considering a data-lake is that the original raw (meta)data is still available for custom applications. A data-lake differs from a data warehouse in that it stores not only structured data but also semi-structured and unstructured data.
To make a data-lake navigable and prevent it from becoming a data swamp, a metadata catalog indexes the stored data by its metadata, as illustrated in the “Functional architecture of a data lake” (Laurent and Madera 2020, Ch. 1). Such metadata management in data lakes is furthermore characterized in Sawadogo et al. (2019) and references therein. Additionally, see Boukraa, Bala, and Rizzi (2024) for a review of metadata management in data lakes.
This work discusses a special case of the data-lake: the metadata-lake [^2] One of the first mentions of the term “metadata-lake” is in Sankar (2021)., for which the admissible data is constrained to metadata encoded in plain-text. Given a setting where such a metadata-lake ingests metadata from various distributed data sources, it induces a virtual data-lake (see Section 4), due to the unified access of distributed data sources, while being its own metadata catalog, cf. Serra (2024), Ch. 6. A metadata-lake can also be seen as a kind of data catalog, as it represents an inventory of available datasets. However, a data catalog product typically builds a database from metadata fields in linked databases (such as column descriptions), while the metadata-lake incorporates metadata from records in linked databases (such as bibliographic metadata). Another related metadata system is the data map (Rooney et al. 2021), which compared to the metadata-lake differs by focusing on tracking metadata changes, whereas the metadata-lake provides an aggregation of current metadata records.
@@
Beyond formally defining a metadata-lake, also the (open-source) implementation of the metadata-lake “DatAasee [^3] The Aasee is an artificial lake in Münster, Germany. (https://github.com/ulbmuenster/dataasee) is described. DatAasee is designed to centralize and unify bibliographic data and metadata of research data and aimed at universities and research institutes, with explicit application in the University of Münster as part of the activities of the University and State Library of Münster.
This work is organized as follows: In Section 2 the potential use-cases for a metadata-lake are detailed. Section 3 reviews the types of metadata and derives a data model for metadata. Then in Section 4 an abstract definition for a metadata-lake is constructed, while in Section 5 the metadata-lake implementation DatAasee is described. Finally, Section 6 briefly evaluates DatAasee in terms of existing metadata system features and presents alternative criteria.
@@ 2 Use-Cases
Before describing the metadata-lake, five potential use-cases and applications for such a service are summarized. All of these were motive for investigating the metadata-lake concept and implementing the DatAasee metadata-lake.
@@ 2.1 Centralized Metadata
First and foremost, the purpose of the proposed metadata-lake is to centralize metadata from multiple sources, to provide a single point of access in a uniform format for university-wide (or company-wide) metadata and thus ultimately data, cf. Anderson (2022). This means distributed data is made available through its metadata by data-pipelining into a central repository, which is the central theme in data engineering (Reis and Housley 2020).
@@ 2.4 FAIR-Compliant Repositories
The FAIR (Findable, Accessible, Interoperable, Reusable) principles (Wilkinson 2016) formalize good scientific practices discipline-agnostically. However, even if particular research is FAIR, research in general can only be FAIR if it, and particularly its metadata, is managed in a repository supporting FAIRness.
@@ 3.1 Intra-Object Metadata
  • Descriptive Metadata, also known as business metadata, refers to information about the contents of a dataset, such as the bibliographic metadata mentioned above as well as license and rights information, and is denoted subsequently by $m_{\text{descriptive}}$.
@@
  • Administrative Metadata determines reading eligibility for the metadata itself as well as the underlying dataset, and is denoted by $m_{\text{administrative}}$.
  • Social Metadata (Yassin 2022) refers to metadata related to discoverability and usage, such as keywords from controlled and uncontrolled vocabularies or number of views, and is denoted by $m_{\text{social}}$.
@@
Among these none is generally elevated, yet $m_{\text{raw}}$ is defining for this work. Assuming metadata management using intra-object metadata without $m_{\text{raw}}$ yields a data warehouse, while using intra-object metadata of exclusively $m_{\text{raw}}$ results in an unmanaged data-lake.
@@ 3.2 Inter-Object Metadata
  • Similarity relates datasets which are, for example, the same work, manifestation, expression, or item using the terminology of FRBR (Functional Requirements for Bibliographic Records 2009); and is symbolized by $m_{\text{similarity}}$.
@@ 3.3 Global Metadata
  • Context information, such as database comments, for instance in SQL via COMMENT ON DATABASE.
@@ 4 Metadata-Lake
Now, given the aim of this work, a metadata-lake cannot just be defined as a set of metadata records; beyond holding a record identifier and the original raw metadata, an internal metadata-lake format needs additional fields to provide a metadata catalog and fulfill the previously presented use-cases.
This format is derived using the metadata classification from Section 3. Overall, the metadata-lake is made up of a hierarchy of tuples:
@@
Intra-object attributes conform to Definition 1, since one can implicitly insert the word “is” between a key and its value and a statement is created of the form “$K_{i}$ is $V_{i}$”. In practical technical terms intra-object metadata is a special case of a key-value document, typical for document databases.
@@
If the labels conform to the (usual) constraint of being a predicate, then also by “$A_{i}$ $R_{i}$ $B_{i}$” statements are given, cf. Semantic Triple (Wikipedia contributors 2024), and the inter-object relations fit the metadata Definition 1.
@@
With respect to Sawadogo et al. (2019), a metadata-lake can then be formally defined by:
@@
A metadata-lake is an ordered triple $(G,T,S)$, where $G$ is a metadata graph, $T$ is a set of transformations, and $S$ is a set of sources, such that $G:=\bigcup_{i=1}^{|S|}T_{f(i)}(S_{i})$ for some suitable map $f:\{1\dots|S|\}\to\{1\dots|T|\}$, with ||$\cdot$|| ⋅ | denoting a set’s cardinality.
This definition is closely related to the data-lake model in Diamantini et al. (2018), Sec. 3, with the distinction that the sources $S$ are metadata sources instead of data sources ($D$), and the addition of transformations $T$.
Also, this abstract definition of a metadata-lake induces already all components for a practical implementation. First, some kind of database [^6] The downstream access to the stored metadata can be realized as a semantic layer. to hold the metadata graph $G$, second, a processor providing transformations $T$ from source formats to the metadata-lake’s internal format, and third, a configuration holding an encoding of the sources $S$, so that the contents of $S$ can be transformed by $T$ and included into $G$. Figure 1 illustrates the full hierarchy of the metadata-lake.
@@ 5 DatAasee
Figure 2: ‘DatAasee‘ metadata-lake.
@@
  • Components are isolated in containers and implemented as declarative as possible.
  • Incoming requests are handled via an HTTP-API and in JSON encoding.
  • State is minimal.
As stated earlier, the primary goal of the metadata-lake is centralizing metadata access, and since the metadata-lake is a data-lake and at the same time also a metadata catalog, the “Hub and Spoke Metadata Architecture” from Laurent and Madera (2020), Sec. 4.8.2 is abstractly employed.
In a top-down approach the metadata-lake architecture is subsequently staked out in ever more detail. For an overview see Figure 2 and Figure 3.
@@ 5.1 Data Architecture
Figure 3: ‘DatAasee‘ outward architecture.
The most foundational decision about building a metadata management service is its data architecture. Fundamentally, the data architecture is fixed already as data-lake, but only to the degree of Definition 2. Implementing a data-lake requires more architectural considerations (Hai et al. 2023).
@@ 5.1.1 Data-Lake Architecture
A process pattern that is closely related to data-lakes is ELT (Extract-Load-Transform), which summarizes a process that extracts data from a source system, loads it as-is into the destination system, and transforms it on-demand. On the other hand, the corresponding process pattern in data warehouses is ETL (Extract-Transform-Load), where data is extracted from a source system, transformed according to the destination system’s schema, and loaded into it. The metadata-lake stores raw metadata like a data-lake, as well as transformed data like a data warehouse due to its metadata catalog aspect. Hence, the EtLT (Extract-transform-Load-Transform) process patterns from Densmore (2021) is used. This means the raw metadata is partially transformed to match the metadata-lake’s catalog schema, while retaining the original raw metadata which is then transformed to its (external) application upon request.
@@ 5.2 Software Architecture
Figure 4: ‘DatAasee‘ inward architecture.
@@ 5.2.1 Data-Tier
A database stores all types of metadata summarized in Section 3. Here the inter-object relations suggest support of a graph model (Ravat and Zhao 2019), while the intra-object attributes imply a document (or relational) model. A property-graph database matches both needs by using documents as nodes and edges. Thus, the proposed metadata model translates as follows: The intra-object attributes correspond to the nodes (vertices), and the inter-object relations are realized by edges, cf. entity-relationship model.
Practically, the DBMS ArcadeDB (Garulli and others, n.d.) is employed in DatAasee. ArcadeDB is a multi-model NoSQL database featuring a hierarchy of models: An overarching graph model with a document model for nodes and edges, and a key-value model for document contents, which results in a property-graph.
A specialty of ArcadeDB is its polyglotism: While the native query language is a dialect of SQL, Gremlin and Cypher are also supported. So, the commonplace SQL eases entry, while domain-specific graph query-languages are available, too.
@@ 5.2.1.1 Data Model
The data model of the DatAasee data-tier follows the metadata-lake schema proposed in Section 3 and Section 4: Intra-object metadata is stored in a vertex, and inter-object metadata is encoded by edges. The vertices consist of descriptive, technical, processual, social, and raw metadata, where the descriptive metadata key-value pairs are derived from the DataCite 4.5 metadata kernel (DataCite Metadata Working Group 2024). Administrative metadata is omitted, since all metadata records are unrestrictedly available to all users. In terms of a relational database, the vertex schema (intra-object attributes) corresponds to a wide table with a star schema for enumerable dimensions. Edge types are a subset of DataCite item relations and the FRBR entity relations. For details about the database schema, see the Appendix.
In the metadata-lake, the two mandatory intra-object attributes are the record identifier and raw metadata. While the former is computed as a hash from a subset of descriptive and processual metadata which is then base64url encoded, the latter is stored as a plain string, since only textual metadata was assumed.
@@ 5.2.2 Application-Tier
Basically, the domain logic consists of two components, the import of metadata on the one hand, and the provisioning of it on the other. The import from source systems includes the request-driven transport, parsing, (partial) transformation, data-cleaning, metadata quality evaluation and storing in the data-tier. Provisioning happens by listening for presentation-tier requests, translating those to data-tier queries, and retranslating the data-tier’s response to the domain response. Requests for a single record by its identifier or search by facets, full-text or custom query in a compatible query-language are supported.
@@ 5.2.2.1 Data Pipelines
The source and transformation components of the metadata-lake, with itself as implicit destination, form data pipelines. A data pipeline needs to be specialized by a configuration to be able to admit updates or additions during operation. Four (metadata) items are minimally needed to specify a data source and transformation for DatAasee:
@@
  • Format: then describes how the content is arranged within the encoding and its syntax, i.e.: DataCite.
Location and protocol are required to extract the metadata from its source, while encoding and format are essential to transform the metadata into the native metadata-lake schema. Potentially, also credentials (i.e.: username/password, access-/secret-key) maybe necessary to gain access to a source. In addition, DatAasee requires a data-steward to be named for every ingested source.
Currently, DatAasee supports the OAI-PMH and S3 protocols, as well as DataCite, DublinCore, MARC, and MODS formats [^7] See openarchives.org/pmh, docs.aws.amazon.com/s3, dublincore.org, loc.gov/standards/mods, and loc.gov/standards/marcxml respectively. Encoding-wise XML is implicitly assumed, but is principally extensible to JSON, CSV or others.

### v3

@@ Abstract
Metadata management for distributed data sources is a long-standing but ever-growing problem. To counter this challenge in a research-data and library-oriented setting, this work constructs a data architecture, derived from the data-lake: the metadata-lake. A proof-of-concept implementation of this proposed metadata aggregator is presented, too, and also evaluated.
@@ 1 Introduction
In the shadows of data evolving to big data, metadata became big metadata. And due to metadata’s idiosyncrasy of describing other data, metadata management for big data warrants its own specialized discussion. In particular the inert dependency of metadata on its data needs to be accounted for, and leads to the driving conjecture behind this work and its object of study: existence, detail and accuracy of metadata determines discoverability, use and value of its associated data. To this end, a system for metadata management, which centralizes metadata from distributed data sources, is defined and its implementation described in the following.
This work focuses expressly on university libraries; and from a library perspective, the overall concept corresponds to a union catalog (Caplan 2003, Ch. 4) of publications and research data. Considering then that data operations (DataOps) is intertwined with metadata management (Reis and Housley 2022, Ch. 6; Strengholt 2023, Ch. 9), operationalizing research data founds upon enabling libraries to manage distributed research data’s metadata.
A starting point is the abstract data architecture: A data-lake is a central repository for the storage of (raw) data, typically together with at least some facility to ingest data (Mathis 2017; Serra 2024). The rationale for considering a data-lake is that the original raw (meta)data is still available for custom applications. A data-lake differs from a data warehouse in that it stores not only structured data but also semi-structured and unstructured data.
To make a data-lake navigable and prevent it from becoming a data swamp, a metadata catalog indexes the stored data by its metadata, as illustrated in the “Functional architecture of a data lake” (Laurent and Madera 2020, Ch. 1). Such metadata management in data lakes is furthermore characterized in Sawadogo et al. (2019) and references therein. Additionally, see Boukraa, Bala, and Rizzi (2024) for a review of metadata management in data-lakes.
This work discusses a special case of the data-lake: the metadata-lake [^2] One of the first mentions of the term “metadata-lake” is in Sankar (2021)., for which the admissible data is constrained to metadata encoded in plain-text. Given a setting where such a metadata-lake ingests metadata from various distributed data sources, it induces a virtual data-lake (see Section 4), due to the unified access of distributed data sources, while being its own metadata catalog, cf. Serra (2024), Ch. 6. A metadata-lake can also be seen as a data catalog, as it represents an inventory of available datasets. However, a data catalog typically builds a database from (global) metadata fields in linked databases (such as column descriptions), while the metadata-lake incorporates (descriptive) metadata from records in linked databases (such as bibliographic metadata). Another related metadata system is the data map (Rooney et al. 2021), which compared to the metadata-lake differs by focusing on tracking metadata changes, whereas the metadata-lake provides an aggregation of current metadata records.
@@
Beyond formally defining a metadata-lake, also the (open-source) implementation of the metadata-lake “DatAasee [^3] The Aasee is an artificial lake in Münster, Germany. (https://github.com/ulbmuenster/dataasee) is described. DatAasee is designed to centralize and unify bibliographic data and metadata of research data and aimed at universities and research institutes, with explicit application at the University and State Library of Münster.
Overall, this work is organized as follows: In Section 2 the potential use-cases for a metadata-lake are detailed. Section 3 reviews the types of metadata and derives a data model for metadata. Then in Section 4 an abstract definition for a metadata-lake is constructed, while in Section 5 the metadata-lake implementation DatAasee is described. Finally, Section 6 briefly evaluates DatAasee in terms of existing metadata system features and also proposes alternative criteria.
@@ 2 Use-Cases
Before describing the metadata-lake, six potential use-cases and applications for such a service are summarized. All of these were motive for investigating the metadata-lake concept and implementing the DatAasee metadata-lake.
@@ 2.1 Centralized Metadata
First and foremost, the purpose of the proposed metadata-lake is to centralize metadata from multiple sources, to provide a single point of access in a uniform format for university-wide (or company-wide) metadata and thus ultimately data, cf. Anderson (2022). This means distributed data is made available through its metadata by data-pipelining into a central repository, which is the central theme in data engineering (Reis and Housley 2022).
@@ 2.4 FAIR-Compliant Repositories
The FAIR (Findable, Accessible, Interoperable, Reusable) principles (Wilkinson 2016) formalize good scientific practices discipline-agnostically. However, even if a particular research artifact is FAIR, research in general can only be FAIR if it, and particularly its metadata, is managed in a repository supporting FAIRness.
@@
2.5 Dataspace
A dataspace is a data management ansatz based on a set of data sources and relations among their contents, as well as a common interface for accessing the participating data source’s contents agnostic of their format (Curry 2020). By unifying data access to distributed data sources without integrating the data itself, conceptually a dataspace is a graph-based catalog for a virtual data-lake. In this sense, the herein proposed metadata-lake can instantiate such a dataspace.
@@ 3.1 Intra-Object Metadata
  • Descriptive Metadata, also known as business metadata, refers to information about the contents of a dataset, such as the bibliographic metadata mentioned above, and is denoted subsequently by $m_{\text{descriptive}}$.
@@
  • Administrative Metadata determines access eligibility for the metadata itself as well as the underlying dataset, and is denoted by $m_{\text{administrative}}$.
  • Social Metadata (Yassin 2022; Strengholt 2023, Ch. 9) refers to metadata related to discoverability and usage, such as keywords from controlled and uncontrolled vocabularies or number of views, and is denoted by $m_{\text{social}}$.
@@
Among these none is generally elevated, yet $m_{\text{raw}}$ is defining for this work. For instance, metadata management using intra-object metadata without $m_{\text{raw}}$ yields a data warehouse, while using intra-object metadata of exclusively $m_{\text{raw}}$ results in an unmanaged data-lake.
@@ 3.2 Inter-Object Metadata
  • Similarity relates datasets which are, the same work, manifestation, expression, or item using the terminology of FRBR (Functional Requirements for Bibliographic Records 2009); and is symbolized by $m_{\text{similarity}}$.
@@ 3.3 Global Metadata
  • Context information, such as database comments, for instance in SQL by COMMENT ON DATABASE (PostgreSQL Global Development Group n. d.).
@@ 4 Metadata-Lake
Now, given the aim of this work, a metadata-lake cannot just be defined as a set of metadata records; beyond holding a record identifier and the original raw metadata, an internal metadata-lake format needs additional fields to provide a metadata catalog and fulfill the previously presented use-cases. Particularly, a field is needed to locate the metadata record’s underlying data if the metadata-lake is used as metadata catalog for a virtual data-lake.
Such a schema is derived using the metadata classification from Section 3. Overall, the metadata-lake is made up of a hierarchy of tuples:
@@
Intra-object attributes conform to Definition 1, since one can implicitly insert the word “is” between a key and its value, and a statement is created of the form $K_{i}$ is $V_{i}$”. In practical technical terms intra-object metadata is a special case of a key-value document, typical for document databases.
@@
If the labels conform to the (usual) constraint of being a predicate, then also by $A_{i}$ $R_{i}$ $B_{i}$” statements are given, cf. Semantic Triple (Wikipedia contributors 2024), and the inter-object relations fit the Metadata Definition 1.
@@
With respect to Sawadogo et al. (2019), a metadata-lake can then be defined by:
@@
A metadata-lake is an ordered triple $(G,T,S)$, where $G$ is a metadata graph, $T$ is a set of transformations, and $S$ is a set of sources, such that $G:=\bigcup_{i=1}^{|S|}T_{f(i)}(S_{i})$ for some suitable map $f:\{1\dots|S|\}\to\{1\dots|T|\}$, with $|\cdot|$ denoting a set’s cardinality.
This definition is closely related to the data-lake model in Diamantini et al. (2018), Sec. 3, with the distinction that the sources $S$ are metadata sources instead of data sources ($D$), and the addition of transformations $T$. Also, the “metalake” from Strengholt (2023), Ch. 9 is a related concept.
This abstract definition of a metadata-lake induces already all components for a practical implementation. First, some kind of database [^6] The downstream access to the stored metadata can be realized as a semantic layer. to hold the metadata graph $G$, second, a processor providing transformations $T$ from source formats to the metadata-lake’s internal format, and third, a configuration holding an encoding of the sources $S$, so that the contents of $S$ can be transformed by $T$ and included into $G$. Figure 1 illustrates the full hierarchy of the metadata-lake.
@@ 5 DatAasee
Figure 2: DatAasee metadata-lake.
@@
  • A functional software architecture with minimal state.
  • Components are isolated in containers.
  • Components are implemented as declarative as possible.
  • Interaction is exclusively realized through an HTTP-API.
  • All API communication is in JSON encoding.
As stated earlier, the primary goal of the metadata-lake is centralizing metadata access, and since the metadata-lake is a data-lake and at the same time also a metadata catalog, the “Hub and Spoke Metadata Architecture” from Laurent and Madera (2020), Sec. 4.8.2 is abstractly employed. Its weakness of a customized standard meta model is mitigated by only partially parsing ingested metadata records and thus reducing the extent of overall load.
In a top-down approach the metadata-lake architecture is subsequently staked out in ever more detail. For an overview of the practical data model (see also Section 5.2.1.1) and the application of the metadata-lake, see Figure 2 and Figure 3.
Figure 3: DatAasee outward architecture.
@@ 5.1 Data Architecture
The most foundational decision about building a (metadata management) service is its data architecture. Fundamentally, the data architecture is fixed already as data-lake, but only to the degree of Definition 2. Implementing a data-lake requires additional practical architectural considerations (Hai et al. 2023).
@@ 5.1.1 Data-Lake Architecture
A process pattern that is closely related to data-lakes is ELT (Extract-Load-Transform), which summarizes a process that extracts data from a source system, loads it as-is into the destination system, and transforms it on-demand. On the other hand, the corresponding process pattern in data warehouses is ETL (Extract-Transform-Load), where data is extracted from a source system, transformed according to the destination system’s schema, and loaded into it. The metadata-lake stores raw metadata like a data-lake, as well as transformed data like a data warehouse due to its metadata catalog aspect. Hence, the EtLT (Extract-transform-Load-Transform) process patterns from Densmore (2021) is effectively used. This means the raw metadata is partially transformed to match the metadata-lake’s catalog schema, while retaining the original raw metadata which is then transformed to its (external) application upon request.
@@ 5.2 Software Architecture
Figure 4: DatAasee inward architecture.
@@ 5.2.1 Data-Tier
Similar to Prabhune et al. (2016), DatAasee is based on a NoSQL database system, which stores all types of metadata summarized in Section 3. Here the inter-object relations suggest support of a graph model (Ravat and Zhao 2019), while the intra-object attributes imply a document (or relational) model. A property-graph database matches both needs by using documents as nodes and edges. Thus, the proposed metadata model translates as follows: The intra-object attributes correspond to the nodes (vertices), and the inter-object relations are realized by edges, cf. entity-relationship model.
Practically, the DBMS ArcadeDB (Garulli et al. n. d.) is employed in DatAasee. ArcadeDB is a multi-model NoSQL database featuring a hierarchy of models: An overarching graph model with a document model for nodes and edges, and a key-value model for document contents, resulting in a property-graph.
A specialty of ArcadeDB is its polyglotism: While the native query language is a dialect of SQL, Gremlin and Cypher are also supported. So, the commonplace SQL eases entry, while specific graph query-languages are available, too.
@@ 5.2.1.1 Data Model
The data model of the DatAasee data-tier follows the metadata-lake schema proposed in Section 3 and Section 4: Intra-object metadata is stored in a vertex, and inter-object metadata is encoded by edges. The vertices consist of descriptive, technical, processual, social, and raw metadata, where the descriptive metadata key-value pairs are derived from the DataCite 4.5 metadata kernel (DataCite Metadata Working Group 2024). Furthermore, administrative metadata is omitted, since all metadata records are unrestrictedly available to all users for reading (license and rights fields are kept as descriptive metadata).
In terms of a relational database, the metadata vertex schema (intra-object attributes) would correspond to a wide table with a star schema for enumerable dimensions. Edge types are a subset of the DataCite item relations together with a subset of the OpenWEMI (Dublin Core Metadata Initiative 2024) entity relations, of which the latter encode the FRBR. For details about the database schema, see the Appendix.
In the metadata-lake, the two mandatory intra-object attributes are the record identifier and raw metadata. While the former is computed as a xxh64 hash from a subset of descriptive and processual metadata and then base64url encoded, the latter is stored as a plain string, since only textual metadata was assumed.
@@ 5.2.2 Application-Tier
Basically, the domain logic consists of two components, the import of metadata on the one hand, and the provisioning of it on the other. The import from source systems includes the request-driven transport, parsing, (partial) transformation, data-cleaning, metadata quality evaluation and storing in the data-tier. Provisioning happens by listening for presentation-tier requests, translating those to data-tier queries, and re-translating the results to a domain response, akin to a semantic layer. Requests for a single record by its identifier or search by filter, full-text or custom query in a compatible query-language are supported.
@@ 5.2.2.1 Data Pipelines
The source and transformation components of the metadata-lake, with itself as implicit destination, form data pipelines. A data pipeline needs to be specialized by a configuration to be able to admit updates or additions during operation. Technically, four (metadata) items are minimally needed to specify a data source and transformation for DatAasee:
@@
  • Format: then describes how the content is arranged, i.e.: DataCite.
Location and protocol are required to extract the metadata from its source, while encoding and format are essential to (partially) transform the metadata into the native metadata-lake schema. Potentially, also credentials (i.e.: username/password, access-/secret-key) maybe necessary to gain access to a source. Non-technically, DatAasee requires a data-steward to be named for every ingested source, as well as the rights (license) under which the metadata of the source are ingested. Beyond these, an options field is available to pass protocol specific configuration, i.e. selective harvesting for OAI-PMH.
Currently, DatAasee supports the OAI-PMH, S3, and GET protocols, as well as DataCite, DublinCore, LIDO, MARC, and MODS formats [^7] See openarchives.org/pmh, docs.aws.amazon.com/s3, httpwg.org/specs/rfc9110, dublincore.org, lido-schema.org, loc.gov/standards/marcxml, and loc.gov/standards/mods. Encoding-wise XML is implicitly assumed, but is principally extensible to JSON, CSV or others. Additionally, the contents of another DatAasee instance can be ingested.
For exporting metadata, DatAasee provides its native format which is derived from DataCite, as well as actual DataCite, and BibJSON (MacGillivray and Pitman n. d.) - a JSON representation of BibTEX.